OpenStudy (anonymous):

How do I find the curve of intersection of this two surfaces z = 1-x^2 and y = 1-x

OpenStudy (nottim):

Curve? Not point?

OpenStudy (nottim):

Well, rearrange them so they look like their more-known forms. Also, try making agraph of these 2.

OpenStudy (anonymous):

yes the intersection of a plane and the other surface hehe.

OpenStudy (nottim):

z=x^2-1 and y=x-1

OpenStudy (nottim):

WAT. I CANT DO THT.

OpenStudy (nottim):

THATS VECTORS>...

OpenStudy (nottim):

Sorry....

OpenStudy (nottim):

i better do my vecotrs work tho

OpenStudy (nottim):

other people, be my savior!

OpenStudy (lgbasallote):

i was thinking systems...but to do it a third equation is needed...is integration applicable?

OpenStudy (anonymous):

Well I don't want to find a point. I want to find a curve. so I just need a pair of equations. But i can figure out a way of getting the equation of the curve.

OpenStudy (lgbasallote):

which math lesson is this? maybe i'll have an idea on how to solve it from there.

OpenStudy (anonymous):

Well I was studying triple integrals.

OpenStudy (lgbasallote):

ohh...then this really has something to do with integration...which i am weak at :p sorry

OpenStudy (anonymous):

Well the integration part is clear, the problem is to get the cuve of intersection of those surfaces =/

OpenStudy (amistre64):

z = 1-x^2 y = 1-x -> x = 1-y z = 1-(1-y)^2 z = 1 - 1+2y-y^2 z = 2y-y^2

OpenStudy (anonymous):

Thanks amistre that is one of the results i got. But this one confuse me too:\[z=1-x^2=(1+x)(1-x)=y(1+x)\]

OpenStudy (anonymous):

I graphed \[z=y(1+x)\]and I got a weird surface. That confused me.

OpenStudy (anonymous):

and the equation \[z=2y-y^2\] it's a surface too. I don't know how does a curve's equation looks like.

OpenStudy (amistre64):

the "curve" is the intersection of the plane and the cylindar; but id have to plot them to see it better

OpenStudy (anonymous):

Here it is:

OpenStudy (amistre64):

yep, thats the setup i had on my paper :)

OpenStudy (anonymous):

So the idea I have on my mind is that the intersection between those two surfaces is a space curve. but none of the equations I have got is a space curve.

OpenStudy (anonymous):

I can interpret z = 2y - y^2 and x = 0. as a curve. But I'm not sure.

OpenStudy (amistre64):

the projection onto the zy plane is 2y-y^2 i wonder if instead of trying to solve for xyz we introduce a x(t), y(t) and z(t)

OpenStudy (amistre64):

x=1-t y = t z=2t-t^2

OpenStudy (anonymous):

I did that and I get this:

OpenStudy (anonymous):

but with x = t

OpenStudy (amistre64):

http://www.wolframalpha.com/input/?i=polar+r%3D%3C1-t%2Ct%2C2t-t%5E2%3E this aint it but its a cool mistake nonetheless

OpenStudy (anonymous):

ahh that it's in polar coordinates. How do I do to specify cartesian coordinates?

OpenStudy (amistre64):

i dunno yet, wolfram hates me

OpenStudy (anonymous):

Ohh that looks better

OpenStudy (amistre64):

that looks like it :)

OpenStudy (anonymous):

Well thank you amistre64!

OpenStudy (amistre64):

yw